OpenStudy (anonymous):

what is the inverse of f(x) = x^2 - 16. Im taking algebra 2 over the summer and im really lost now taking the pretest...

OpenStudy (anonymous):

Plato?

OpenStudy (anonymous):

y=x^2-16 x->y and y->x x=y^2-16 x+16=y^2 \[y=\sqrt{16+x}\]